### v1

# Improving Autoformalization Using Direct Dependency Retrieval

Abstract
The convergence of deep learning and formal mathematics has spurred research in formal verification. Statement autoformalization, a crucial first step in this process, aims to translate informal descriptions into machine-verifiable representations but remains a significant challenge. The core difficulty lies in the fact that existing methods often suffer from a lack of contextual awareness, leading to hallucination of formal definitions and theorems. Furthermore, current retrieval-augmented approaches exhibit poor precision and recall for formal library dependency retrieval, and lack the scalability to effectively leverage ever-growing public datasets. To bridge this gap, we propose a novel retrieval-augmented framework based on DDR (Direct Dependency Retrieval) for statement autoformalization. Our DDR method directly generates candidate library dependencies from natural language mathematical descriptions and subsequently verifies their existence within the formal library via an efficient suffix array check. Leveraging this efficient search mechanism, we constructed a dependency retrieval dataset of over 500,000 samples and fine-tuned a high-precision DDR model. Experimental results demonstrate that our DDR model significantly outperforms SOTA methods in both retrieval precision and recall. Consequently, an autoformalizer equipped with DDR shows consistent performance advantages in both single-attempt accuracy and multi-attempt stability compared to models using traditional selection-based RAG methods.
Machine Learning, ICML
1 Introduction
Theorem provers, such as Lean (moura2021lean), Coq (bertot2013interactive), and Isabelle (nipkow2002isabelle), are symbolic systems designed for formal verification whose soundness and completeness are proven in theory. Grounded in specific foundational logics, these systems employ proprietary formal languages to encode mathematical concepts, statements, and proofs. The logical kernel within each system can then verify the validity of these formalized objects with complete rigor (nawaz2019survey). However, this process eschews the flexibility of natural language, demanding that users not only master the strict formal syntax but also possess a profound understanding of the semantic mapping between mathematical ideas and their formal representations. Consequently, the effective utilization of these theorem provers presents a formidable challenge to researchers (li2024survey).
Statement autoformalization refers to the process of automatically translating mathematical statements from natural language into a formal language comprehensible to theorem provers (liu2025rethinking). As a foundational step in the automated theorem proving, the correctness of its output is paramount. It determines whether the proof target of a subsequent prover is precisely aligned with the intent of the original statement. Beyond this core function, statement autoformalization exhibits significant potential for broader applications and innovation. These applications include synthesizing training data for neural theorem provers (lin2025goedel), automated program verification (lin2024fvel), curating and filtering informal reasoning datasets (zhang2025deeptheorem), and evaluating and enhancing the informal reasoning capabilities of large language models (e.g., through techniques like rejection sampling (zhou2024don)).
Current mainstream research in autoformalization primarily encompasses three direction. First, constructing larger and more challenging datasets for SFT (Supervised Fine-Tuning) (cao2025informal; peng2025criticlean; yu2025formalmath; miranda2025veribench; kimina_prover_2025; ying2024lean; zheng2021minif2f; azerbayev2023proofnet); second, encoding and indexing the libraries of theorem provers to support retrieval (liu2025rethinking; yang2023leandojo; wang2025aria); and third, employing methods like ICL (In-Context Learning) to enable LLMs (Large Language Models) to directly translate informal statements into formal code (wu2022autoformalization; zhang2024consistent).
However, these approaches face two critical limitations. First, relying solely on generative models to directly synthesize formal code often overlooks the context of existing definitions and lemmas within the formal library (yang2023leandojo; zhang2024consistent). This lack of contextual awareness is a primary reason for hallucinations in the model’s use of identifiers and construction of syntax . For complex theorems, effectively leveraging relevant conclusions from the library is crucial for reducing formalization difficulty. Without this capability, LLMs are forced to locally reinvent functionally equivalent but more complex expressions or to selectively ignore key premises they cannot handle, which readily introduces errors. Similiar results have been reported in wu2022autoformalization and yu2025formalmath, pointing out that definitional mismatches between informal mathematics and the formal library, as well as the inappropriate usage of tactics and theorems are primary causes of formalization failures. Figure 1 (a) intuitively illustrates this issue: unable to access the predefined Nat.factorization in the library, the model attempts to construct an equivalent operator locally, resulting in erroneous code. Similarly, the results in Table 3, which show a significant performance decline for all ICL-based models as task difficulty increases, further underscore the severity of this problem. Second, the retrieval strategies for library dependencies in existing methods are often rudimentary, yielding results with weak semantic relevance to the informal statement (azerbayev2023proofnet; yang2023leandojo). The libraries of mainstream theorem provers are vast and continually growing; for instance, Lean 4’s mathlib contains over 240,000 entries, including many with similar descriptions and usage. Simple retrieval methods struggle to achieve both high efficiency and precision. This challenge is confirmed by the experiments in liu2025rethinking as well as our own results presented in Table 2.
To address the aforementioned challenges, we propose a novel dependency retrieval method DDR, and the corresponding framework for statement autoformalization. In this context, dependency retrieval aims to identify relevant formal objects, such as definitions and theorems, from a theorem prover’s library based on an informal mathematical statement. Unlike traditional selection-based retrieval methods, DDR leverages the contextual understanding of LLMs to directly generate potential formal dependencies. To empower the LLM with this capability, we have designed a suffix array-based methodology for both data generation and dependency verification. Specifically, this process begins by constructing an efficient database from the theorem prover’s library using a suffix array. Subsequently, leveraging existing informal-formal statement pairs from public datasets, we extract candidate dependencies from the formal code and verify them against the database via the efficient binary search. This produces high-quality labels for fine-tuning the DDR model. The same verification mechanism is reused at inference time to validate the dependencies generated by the DDR model. Given the ever-increasing size of modern theorem prover libraries, we place special emphasis on the efficiency of the verification step in our RAG pipeline. Let $N$ denote the volume of the library, the time complexities for database construction and verification in our proposed method are $\mathcal{O}(N)$ and $\mathcal{O}(\log N)$ respectively, ensuring the scalability of our approach to libraries of any size. By applying the above process to the FineLeanCorpus dataset, we successfully extracted all Mathlib dependencies for over 500,000 samples, reducing the processing time from approximately 2.5 days to under 2 hours. This achievement has enabled the construction of a large-scale dataset for fine-tuning and paves the way for the online deployment of RAG services. Experimental results demonstrate that DDR significantly improves both recall and precision in dependency retrieval compared to traditional methods. Consequently, an autoformalizer equipped with DDR exhibits consistent advantages in both single-attempt accuracy and multi-attempt stability over autoformalizers using conventional RAG approaches.
Our main contributions are summarized as follows:
  1. 1.
    We propose a library dependency verification process based on suffix array. With its $\mathcal{O}(\log N)$ verification time complexity, this highly efficient and scalable method provides critical support for deploying online RAG services and constructing large-scale fine-tuning datasets.
  2. 2.
    We introduce an innovative RAG method based on a generation-plus-verification paradigm. The core component, DDR (Direct Dependency Retrieval), unleashes the contextual awareness of LLMs, enabling them to directly generate potential library dependencies from informal input, enhancing the precision and recall of retrieval results.
  3. 3.
    We conducted comprehensive experiments on statement autoformalization. The results validate that DDR consistently outperforms baseline models that rely on traditional retrieval methods.
  4. 4.
    We have constructed and open-sourced a library dependency dataset with over 500,000 samples. Each sample consists of an informal statement and its corresponding formal library dependencies, suitable for fine-tuning LLMs. The methodology for constructing this dataset is generalizable and can be applied to other formal corpora, fostering more diverse research in statement autoformalization.
Figure 1: Comparison of two paradigms for formal dependency retrieval. (a) Select-based retrieval using embedding cosine similarity: both the input statement and library items are encoded into embeddings, and the top-k items with highest cosine similarity are retrieved as dependencies. (b) DDR (Direct Dependency Retrieval): an LLM generates potential dependencies based on context, followed by a suffix-array-based verification step. DDR achieves higher retrieval accuracy and scalability with a more direct and efficient workflow.
2 Related Work
Formal mathematics aims to translate mathematical statements and proofs into machine-verifiable formal languages based on rigorous logical calculus. In recent years, with the advancement of deep learning, this field has become an active area of research, primarily encompassing two core branches: autoformalization (murphy2024autoformalizing; liu2025rethinking; poiroux2024improving; azerbayev2023proofnet; wang2020exploration), and automated theorem proving (lin2025goedel; kimina_prover_2025; chen2025seed; xin2024deepseek; xin2024deepseekv15; ren2025deepseek; liu2025proofaug).
Automated theorem proving is dedicated to automatically generating proof scripts for given formal goals. The central challenge lies in designing efficient proof search strategies (liu2025proofaug). Current cutting-edge research in this direction focuses on combining LLMs with reinforcement learning to emulate the heuristic reasoning processes of human mathematicians in exploring the proof space (chen2025seed), with the future prospect of assisting in the discovery of new theorems. Prominent examples of such systems include Seed-Prover (chen2025seed) and DeepSeek-Prover (xin2024deepseek; xin2024deepseekv15; ren2025deepseek). In contrast, autoformalization targets the automatic translation of informal mathematical content, typically expressed in natural language, into a formal language. Research in this area concentrates on establishing reliable evaluation methodologies to assess the semantic fidelity of the translation (liu2025rethinking; poiroux2024improving) and on developing higher-performance autoformalization models (liu2025rethinking; poiroux2024improving). This paper focuses on the task of statement autoformalization. We propose a method to mitigate model hallucination and enhance model performance by offering better dependency retrieval results in the RAG framework.
Progress in autoformalization research has led to the creation of large-scale informal-to-formal parallel corpora (mahdavi2025leveraging). These corpora, such as FormalMath (yu2025formalmath), CriticBench (peng2025criticlean), and FineLeanCorpus (peng2025criticlean), have expanded in scale from a few hundred samples to tens of thousands and are primarily generated through rule-based methods and by leveraging the ICL capabilities of LLMs (azerbayev2023proofnet; gao2024omni). However, current autoformalization research has not effectively utilized this data to discover the intricate connection between informal mathematical statements and their required library dependencies. To address this gap, we propose DDR, a method designed to enhance a model’s ability to infer and utilize library dependencies during statement formalization.
Research on evaluating the fidelity of autoformalization has explored various methods, such as manual expert review (wang2018first; peng2025criticlean), LLM review (yang2023leandojo; sidhu2024evaluation), auxiliary metrics like Perplexity (wang2018first) and BLEU (papineni2002bleu; wang2018first), and type checking (poiroux2024improving) within a theorem prover. However, these approaches have inherent limitations in capturing the deep semantic equivalence between informal and formal mathematical statements. They may overlook semantic inconsistencies between the two modalities or produce false negatives by incorrectly rejecting semantically valid formalizations (liu2025rethinking). To overcome this, our work adopts BEq (Bidirectional Extender Definitional Equivalence) as the primary evaluation metric (liu2025rethinking). This choice is motivated by BEq’s high alignment with human expert judgment in assessing semantic equivalence. The core principle of BEq is to define two formal statements as equivalent if they are mutually transformable through a predefined, finite, and verifiable set of transformation rules (aczel2013homotopy).
Figure 2: Workflow of the suffix array auxiliary checking process. All library items are concatenated into a single string with delimiters, followed by suffix array construction. Each pending item is then queried via binary search to determine its match status (exact, partial, or none). This mechanism supports both dataset construction and LLM output verification.
3 Methodology
3.1 Direct Dependency Retrieval
To prevent hallucination during generation, existing dependency retrieval methods select items from a pre-existing library to form their output (yang2023leandojo; liu2025rethinking), as illustrated in Figure 1 (a). However, this selected-based paradigm suffers from two primary limitations. First, training an encoder to produce effective embeddings for every formal object in the library to maximize retrieval performance is a challenge in itself (xiong2020approximate; gao2023retrieval). Second, the method is sensitive to superficial variations in the input: semantically equivalent mathematical statements may yield different results in maximum cosine similarity retrieval due to minor differences in their syntactic structure or notational representation (asl2024robustsentembed). While the first issue can be partially mitigated by techniques such as optimizing the input format for the embedding model, the second remains an challenge. Our empirical results confirm that such embedding-based retrieval methods exhibit low recall and precision. Inaccurate retrieval results don’t contribute to the language model in autoformalization tasks. Figure 1 (a) illustrates this failure mode: in the absence of a usable dependency from the retrieved Mathlib candidates, the model attempts to construct a local function functionally equivalent to factorization, leading to an erroneous output. Furthermore, this approach lacks scalability, as it cannot effectively leverage the growing volume of available datasets to continuously enhance model capabilities.
To address the challenge of LLMs hallucinating non-existent or irrelevant library dependencies during autoformalization, we propose a generation-plus-verification pipeline, with DDR as its core component. DDR is a LLM-based SFT model to operate in an end-to-end fashion, directly extracting potential library dependencies from an input mathematical statement. In contrast to conventional two-stage retrieval methods that require additional encoder model (liu2025rethinking), the design of DDR streamlines the architecture, leading to a simpler, more efficient, and better-performing approach. Our experimental results demonstrate the effectiveness of DDR, showing significant improvements in both recall and precision compared to mainstream embedding-based retrieval methods. Accurate dependency retrieval reduces the complexity of the subsequent formalization task, thereby enhancing the overall performance of the autoformalizer. As illustrated in Figure 1 (b), providing the model with precise dependencies, such as Nat.factorization and Nat.factorial, enables it to more reliably generate the correct formal output, thus increasing the overall success rate. Furthermore, the DDR framework is designed to leverage existing datasets to augment its training corpus, providing a viable path for the continuous improvement of its retrieval accuracy.
The effective application of the DDR method confronts two primary challenges. First, unlike embedding-based approaches that retrieve existing entries from a library, DDR relies on a LLM to directly generate potential dependencies. This generative nature necessitates the verification of its output to filter out non-existent items not present in the formal library (i.e., hallucinations). The complexity of this verification process is exacerbated by name ambiguity: as the model is trained on code, its output often consists of unqualified short names (e.g., sqrt) rather than the fully qualified names of formal objects (e.g., Int.sqrt). Such short names can correspond to multiple entities within the library (e.g., Nat.sqrt or Real.sqrt), posing a challenge for precise matching. Second, training a high-performance DDR model requires specialized datasets. However, existing corpora of formalized mathematics generally lack high-quality paired data that directly links mathematical statements to their required dependencies. Consequently, how to efficiently construct such training data from existing formal code and effectively leverage the ever-growing new corpora presents a critical problem. To systematically address these issues of verification and data construction, we will elaborate on our proposed solutions in the subsequent two subsections.
3.2 Suffix Array Auxiliary
To accurately and efficiently verify the existence of dependency objects generated by LLMs within a formal library, we introduce a checking method based on suffix arrays (liu2024infini). The core idea of this approach is to first concatenate the identifiers of all defined objects in the library, such as pre-proven theorems, into a single long string using a specific delimiter. Subsequently, a suffix array is constructed from this concatenated long string in linear complexity (lee2022deduplicating; karkkainen2006linear). By definition, a suffix array is an array of integers that stores the starting indices of all suffixes of the original string, sorted in lexicographical order (karkkainen2006linear). Figure 2 illustrates the suffix array auxiliary check process with a simplified example, assuming the library contains only two objects and there are three pending items. The sorted nature of the suffix array facilitates efficient matching of these pending items via binary search. Specifically, for each pending item, we can locate the range of all suffixes that have the item as a prefix within the suffix array. This process enables the rapid retrieval of all library objects corresponding to the pending items, thereby completing their existence verification.
The suffix array auxiliary check is a highly efficient method for matching and verification, capable of meeting the real-time requirements of large-scale requests (liu2024infini). For a formal library containing $N$ objects with an average length of $d$, and $M$ pending items with an average length of $s$, conventional string matching methods typically involve brute-force comparison, resulting in a time complexity of $\mathcal{O}\left((s+d)MN\right)$. In contrast, the suffix array auxiliary check operates on a suffix array constructed over the entire library. For each pending item, it utilizes binary search on the suffix array to locate matches. The overall time complexity of this approach is $\mathcal{O}(sM\log N)$. This significant reduction in computational complexity enables real-time responsiveness, thereby making the method suitable for deployment in online application systems (liu2024infini).
3.3 Constructing High-quality Training Dataset
Fine-tuning an accurate and efficient DDR model necessitates a dataset composed of pairs of mathematical statements and their corresponding library dependencies. However, no such public dataset is readily available for this purpose. The primary challenge lies in the difficulty of accurately extracting and efficiently verifying library dependencies from existing formal statements. To address this, we utilize the suffix array auxiliary check to process large-scale autoformalizer datasets and construct the required training data. We chose FineLeanCorpus (peng2025criticlean) as our source, a dataset containing over 500,000 samples that span a wide range of mathematical domains and difficulty levels. After partitioning the data into training and test sets based on problem difficulty, we consolidated all training data and filtered the potential library dependencies. Leveraging the high efficiency of our suffix array-based approach, we processed the entire dataset in under two hours, resulting in a dataset suitable for fine-tuning the DDR model. This resulting dataset has been made open-source at https://huggingface.co/datasets/Palca/FineLeanCorpus_DDR. We anticipate that this contribution will foster further research on RAG-based approaches within the field of formal mathematical proving.
Table 1: Dataset statistics of FineLeanCorpus. Depend Rate denotes the percentage of samples that have explicit mathlib dependencies within each difficulty level. Depend Length denotes the mean number of explicit mathlib dependency items for samples within each difficulty level. Higher Depend Rate and Depend Length values typically indicate more challenging tasks.
Difficulty Num Depend Rate Depend Length 0 835 0.202 395 209 580 838 310.202\,395\,209\,580\,838\,31 0.287 425 149 700 598 80.287\,425\,149\,700\,598\,8 1 120040 0.560 421 526 157 947 40.560\,421\,526\,157\,947\,4 1.074 716 761 079 640 21.074\,716\,761\,079\,640\,2 2 113816 0.785 056 582 554 298 10.785\,056\,582\,554\,298\,1 1.870 738 736 205 805 91.870\,738\,736\,205\,805\,9 3 49841 0.808 330 490 961 256 80.808\,330\,490\,961\,256\,8 2.088 682 008 787 9462.088\,682\,008\,787\,946 4 69882 0.795 927 420 508 857 80.795\,927\,420\,508\,857\,8 1.963 123 551 129 046 11.963\,123\,551\,129\,046\,1 5 67547 0.524 064 725 302 3820.524\,064\,725\,302\,382 1.112 691 903 415 399 71.112\,691\,903\,415\,399\,7 6 41312 0.813 032 532 920 216 90.813\,032\,532\,920\,216\,9 2.251 839 659 178 9312.251\,839\,659\,178\,931 7 36675 0.756 264 485 344 240.756\,264\,485\,344\,24 2.135 460 122 699 386 32.135\,460\,122\,699\,386\,3 8 8682 0.821 354 526 606 772 70.821\,354\,526\,606\,772\,7 2.430 085 233 817 0932.430\,085\,233\,817\,093 9 708 0.830 508 474 576 271 20.830\,508\,474\,576\,271\,2 2.632 768 361 581 9212.632\,768\,361\,581\,921
Table 2: Comparison of our DDR with baseline methods. R@k denotes the topological-informalization retrieval method from liu2025rethinking, where the top-k retrieved results are used as dependency outputs. The symbol $\uparrow$ indicates that our DDR consistently outperforms the baselines.
Difficulty R@5 R@10 DDR Precision Recall F1 $Precision\uparrow Recall\uparrow F1\uparrow$ Diff01 $0.033$ 0.093 333 30.093\,333\,3 0.046 424 963 924 963 930.046\,424\,963\,924\,963\,93 $0.023$ 0.132 916 666 666 666 660.132\,916\,666\,666\,666\,66 0.037 839 452 214 452 2160.037\,839\,452\,214\,452\,216 $0.845$ 0.737 250.737\,25 0.766 428 571 428 571 30.766\,428\,571\,428\,571\,3 Diff23 $0.083$ 0.174 601 190 476 190 470.174\,601\,190\,476\,190\,47 0.104 828 255 078 255 080.104\,828\,255\,078\,255\,08 $0.053$ 0.233 892 857 142 857 120.233\,892\,857\,142\,857\,12 0.081 945 344 687 991 760.081\,945\,344\,687\,991\,76 $0.925$ 0.580 406 746 031 7460.580\,406\,746\,031\,746 0.664 218 253 968 2540.664\,218\,253\,968\,254 Diff45 $0.057$ 0.131 416 666 666 666 650.131\,416\,666\,666\,666\,65 0.074 401 875 901 875 910.074\,401\,875\,901\,875\,91 $0.0375$ 0.173 416 666 666 666 660.173\,416\,666\,666\,666\,66 0.058 731 601 731 601 740.058\,731\,601\,731\,601\,74 $0.91$ 0.683 249 999 999 999 90.683\,249\,999\,999\,999\,9 0.736 361 471 861 471 80.736\,361\,471\,861\,471\,8 Diff67 $0.087$ 0.186 214 285 714 285 750.186\,214\,285\,714\,285\,75 0.110 370 129 870 129 880.110\,370\,129\,870\,129\,88 $0.068$ 0.296 952 380 952 380 90.296\,952\,380\,952\,380\,9 0.104 312 102 113 572 710.104\,312\,102\,113\,572\,71 $0.92$ 0.569 555 555 555 555 60.569\,555\,555\,555\,555\,6 0.651 988 095 238 095 30.651\,988\,095\,238\,095\,3 Diff89 $0.115$ 0.193 134 920 634 920 630.193\,134\,920\,634\,920\,63 0.134 166 527 916 527 940.134\,166\,527\,916\,527\,94 $0.0755$ 0.257 095 238 095 238 10.257\,095\,238\,095\,238\,1 0.110 275 899 109 490 430.110\,275\,899\,109\,490\,43 $0.855$ 0.462 984 126 984 1270.462\,984\,126\,984\,127 0.549 760 442 773 600 70.549\,760\,442\,773\,600\,7
4 Direct Dependency Retrieval-augmented Autoformalization
While the current autoformalization paradigm approach effectively prevents the hallucination of formal objects that do not exist within the library (liu2025rethinking; yang2023leandojo; azerbayev2023proofnet), the accuracy of existing retrieval methods is limited (liu2025rethinking). Furthermore, when the number of retrieved items exceeds the actual number of required dependencies, these superfluous items introduce noise into the autoformalization process. Consequently, such methods typically require prior knowledge of the number of dependencies to mitigate this noise, a prerequisite that is often impractical in application. DDR achieves significantly higher precision and recall in identifying the actual dependencies from Mathlib. It operates without requiring prior knowledge of the number of dependencies and better approximates the distribution of true dependencies, leading to superior generalization capabilities. Comprehensive experimental results validate the effectiveness and superiority of the DDR method.
Table 3: Comparison and ablation results of retrieval-augmented methods. Above the dashed line are the ICL methods; below are the SFT methods. All ICL methods use the same prompt as in liu2025rethinking. MMA denotes the LLM fine-tuned on MMA’s Lean subset. RA denotes the pure autoformalizer method in liu2025rethinking. R@5 and R@10 denote retrieval-augmented methods using top-5 and top-10 retrieved results, respectively. RA+DDR is our main method.
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq DeepSeek-R1 0.502 512 562 814 070 30.502\,512\,562\,814\,070\,3 0.216 080 402 010 050 240.216\,080\,402\,010\,050\,24 $0.255$ $0.17$ 0.277 777 777 777 777 80.277\,777\,777\,777\,777\,8 0.176 767 676 767 676 770.176\,767\,676\,767\,676\,77 0.272 727 272 727 272 70.272\,727\,272\,727\,272\,7 0.191 919 191 919 191 90.191\,919\,191\,919\,191\,9 0.217 171 717 171 717 180.217\,171\,717\,171\,717\,18 0.166 666 666 666 666 660.166\,666\,666\,666\,666\,66 Ling-flash-2.0 $0.635$ $0.575$ $0.365$ $0.33$ $0.425$ $0.41$ $0.305$ $0.3$ $0.17$ $0.16$ Claude-3.5-Sonnet $0.67$ $0.61$ $0.52$ $0.48$ $0.55$ $0.53$ $0.405$ $0.395$ $0.335$ $0.315$ GPT-4o $0.66$ $0.56$ $0.49$ $0.48$ $0.52$ $0.515$ $0.44$ $0.435$ $0.265$ $0.265$ Qwen2.5-Max $0.645$ $0.565$ $0.4$ $0.385$ $0.505$ $0.505$ $0.4$ $0.4$ $0.25$ $0.235$ MMA $0.16$ $0.145$ $0.185$ $0.17$ $0.305$ $0.295$ $0.21$ $0.21$ $0.15$ $0.15$ RA $0.245$ $0.21$ 0.441 340 782 122 9050.441\,340\,782\,122\,905 0.413 407 821 229 050 30.413\,407\,821\,229\,050\,3 0.497 206 703 910 614 50.497\,206\,703\,910\,614\,5 0.491 620 111 731 843 560.491\,620\,111\,731\,843\,56 0.388 888 888 888 888 90.388\,888\,888\,888\,888\,9 0.382 716 049 382 716 030.382\,716\,049\,382\,716\,03 $0.285$ $0.275$ RA+R@5 $0.325$ $0.29$ $0.535$ $0.5$ $0.535$ $0.53$ $0.465$ $0.455$ $0.36$ $0.33$ RA+R@10 $0.335$ $0.3$ $0.48$ $0.45$ $0.545$ $0.535$ $0.475$ $0.455$ $0.31$ $0.31$ RA+DDR $0.36$ $0.315$ $0.56$ $0.52$ $0.60$ $0.59$ $0.54$ $0.51$ $0.41$ $0.38$
4.1 Dataset Overview
To construct and evaluate our proposed DDR method, we used the FineLeanCorpus (peng2025criticlean), a dataset that comprises a vast collection of natural language mathematical statements paired with their corresponding formalizations in Lean 4. To prevent data leakage from potential overlap with other Lean 4 datasets (due to its diverse origins), we split the corpus into training and test sets based on its intrinsic difficulty attribute.
The difficulty attribute is a key metric in FineLeanCorpus designed to quantify the challenge of formalizing informal natural language mathematical statements into Lean 4 statements, rated on an 11-point scale from 0 to 10. In our experimental setup, we consolidated samples with a difficulty of 10 into the difficulty 9 category.
A notable observation is that samples categorized as difficulty 0 and difficulty 1 present a counter-intuitively high formalization challenge. This is primarily because the natural language statements in this category often do not represent abstract descriptions of the problem’s mathematical essence, as illustrated by examples in appendix A. Our subsequent experimental results corroborate this finding: for tasks of difficulty 0 and difficulty 1, the performance of ICL significantly surpasses that of SFT models, a phenomenon that aligns with the analysis.
We applied the suffix algorithm described in Section 3.2 to extract Mathlib dependencies for all statements within each difficulty category. The statistical results are summarized in Table 1. The data reveals a positive correlation between a statement’s general difficulty and its degree of Mathlib dependency: categories of higher difficulty typically exhibit both a higher dependency rate and a greater number of dependencies. This finding aligns with our intuition and preliminary observations of the dataset.
For the experimental setup, we constructed five test sets. Specifically, we randomly sampled 100 statements from each difficulty category and then combined the samples from every two adjacent difficulty categories. This process resulted in five test sets, named Diff01, Diff23, Diff45, Diff67, and Diff89, where Diffij denotes the combined set of samples from difficulty levels i and j, with each test set containing 200 statements. All remaining samples in the dataset were used to fine-tune the DDR model.
4.2 Setup
We employ precision, recall, and F1 score as quantitative metrics to evaluate the performance of our dependency retrieval results. A key challenge in this evaluation arises from the discrepancy between the outputs of DDR, which directly generate dependency names, and the ground-truth labels, which are Mathlib object identifiers extracted from source code. This leads to name ambiguity, as source code often contains unqualified or partially qualified names. To accurately identify true tositives for our evaluation, we introduce a suffix-based matching criterion. Specifically, we represent each identifier as a sequence of dot-separated components. A retrieved identifier is considered to match a ground-truth label if the component sequence of one is a suffix of the other.
To evaluate the semantic equivalence between the formal statements predicted by the model and the ground-truth formal statements, we employ two criteria: TypeC (Type Checking) and BEq (Bidirectional Extended Definitional Equality) (liu2025rethinking). TypeC ensures the syntactic correctness of expressions, whereas BEq is utilized to determine their semantic equivalence. Building upon these criteria, we define the TypeC@k and BEq@k evaluation metrics. These metrics measure the percentage of problems for which at least one of the top-k predicted statements is deemed equivalent to the ground-truth statement under the corresponding criterion.
Figure 3: Comparison of our DDR with baselines. (a)–(c) show results for R@5, R@10, and DDR, respectively. R@k is the topological-informalization retrieval method from liu2025rethinking. Light and wider bars indicate the mean; dark and thinner bars show the standard deviation. Gray: F1; Red: Recall; Green: Precision.
To comprehensively evaluate our proposed method, we established several baseline models and conducted ablation studies. For the ablation studies, we employed the RA (RAutoformalizer) from liu2025rethinking as a pure formalizer baseline, and RA+R@k as an autoformalizer that incorporates the top-k retrieved dependencies using the method in liu2025rethinking. The RA was fine-tuned based on InternLM2-Math-Base-7B (ying2024internlm). Our baselines cover multiple aspects: first, we evaluated a range of state-of-the-art general LLMs via ICL, including DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (anthropic_claude35_sonnet_2024), GPT-4o (achiam2023gpt), and Qwen2.5-Max (qwen2025qwen25technicalreport); second, we included an InternLM2-Math-Base-7B (ying2024internlm) model fine-tuned on the MMA Lean subset (jiang2023multilingual). For our proposed DDR method, we selected Qwen3-32B (yang2025qwen3) as its base model. This choice highlights a core advantage of DDR: its flexibility in selecting a base dependency retrieval model without being restricted to specific embedding models.
During the decoding stage for all experiments, the temperature was set to 0.7, and we generated 8 samples for each problem. The first generated output was used to calculate Metric@1, while all 8 outputs were used for Metric@8. To ensure a fair comparison, the dependency retrieval step for the DDR method was performed only once, and its results were reused for all subsequent Lean 4 statement formalization generations. It is noteworthy that none of the autoformalization models discussed in the main text, including the fine-tuned ones, have undergone SFT on the FineLeanCorpus (peng2025criticlean). Further details on the experimental setup are provided in appendix B.
4.3 Retrieval Results
Table 2 presents the dependency retrieval performance, evaluated on 200 samples across datasets of varying difficulty. The performance is measured by the average Precision, Recall, and F1-Score. Here, R@5 and R@10 denote the results of the RA method when retrieving the top-5 and top-10 most relevant dependencies, respectively.
The R@5 baseline exhibits suboptimal overall performance. As indicated in Table 1, the average dependency length across all difficulty levels is less than 5, which accounts for its low Precision. However, the concurrently low Recall suggests that the retrieved dependencies are inaccurate, leading to poor overall retrieval quality. By increasing the number of retrieved items to 10, R@10 achieves a higher Recall than R@5, but at the cost of a significant drop in Precision, resulting in a lower F1-Score. This phenomenon indicates that the traditional RAG method is not well-suited for autoformalization problems. Furthermore, Figure 3 illustrates the standard deviations of these metrics on 200 samples. The data reveals that the output quality stability, as measured by standard deviation, is comparable among DDR, R@5, and R@10. Nevertheless, DDR consistently and significantly outperforms existing RAG methods across all datasets in terms of the mean of Precision, Recall, and F1-Score. This result highlights the superiority of our proposed DDR approach.
4.4 Autoformalization Results
Table 3 reveals a complex relationship between model performance and dataset characteristics, with several findings that are counter-intuitive.
Figure 4: LLM consistency across multiple attempts. Subplots (a–e) correspond to datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot represents the number of samples ($y$ axis) whose outputs pass the BEq criterion at least $k$ ($x$ axis) times out of 8 attempts. Blue: RA+DDR; Orange: RA+R@10; Green: RAW
Figure 5: LLM efficiency across multiple attempts. Subplots (a–e) correspond to datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot shows the distribution of samples ($y$ axis) by the number of attempts ($x$ axis) required to pass the BEq criterion for the first time over 8 attempts. Blue: RA+DDR; Orange: RA+R@10; Green: RA
On the Diff01 subset, where inputs resemble unstructured natural language, SFT models generally exhibited poor performance. In contrast, ICL models, which are better adapted to general linguistic expressions, performed more effectively. Notably, on the Diff45 subset, which features a low average Depend Rate and Depend Length as shown in Table 1, SFT models demonstrated a performance improvement despite the higher difficulty level. ICL models also showed an anomalous increase in performance on this subset, contradicting the general trend where performance degrades as difficulty increases. For the other subsets, the results consistently show that as the difficulty level ascends, both the average Depend Rate and Depend Length within the dataset increase, corresponding directly with a steady decline in performance across all tested models.
Collectively, these observations establish a strong negative correlation between the performance of autoformalization models and the complexity of library dependency, as quantified by Depend Rate and Depend Length. This finding underscores the necessity of employing techniques such as RAG to aid models in managing such dependencies.
Against this backdrop, our proposed RA+DDR method demonstrates superior performance. As shown in Table 3, our method outperforms all baseline models on the BEq metric across nearly all subsets, including state-of-the-art models with tens of times more parameters. This result strongly indicates that autoformalization is a specialized challenge that requires targeted solutions beyond simple prompt engineering, further validating the value of our in-depth research into RAG.
4.5 Ablation Studies
Figure 4 evaluates the ablation methods from the perspective of multi-attempt reliability. It plots the number of problems for which the model’s generated Lean 4 statements were equivalent to the ground truth for n times (where n ranges from 1 to 8) across 8 independent generation attempts. The results indicate that, compared to the baseline methods RA and RA+R@5, our proposed method, RA+DDR, achieves higher consistency on a larger set of problems. RA+DDR outperforms the baselines, demonstrating the superior stability and reliability of its statement formalization results.
Figure 5 presents a comparison from the perspective of generation efficiency. It illustrates the distribution of problems based on the number of attempts required to achieve the first successful match with the ground truth across 8 independent trials. The experimental results show that the RA+DDR method tends to find the correct statement formalization in fewer attempts. In contrast, the success distribution for the baseline methods, RA and RA+R@5, is skewed towards requiring more trials to achieve the first match. This finding further highlights the efficiency advantage of our proposed method for the autoformalization task.
5 Discussion
DDR departs from traditional retrieval-based methods for formal dependency acquisition by employing an innovative generation-plus-verification paradigm. A key limitation of DDR, however, is its inability to precisely locate the generated items within the library. Specifically, whereas the full retrieval results in methods like liu2025rethinking typically include complete metadata such as the item’s name, code usage, and documentation, DDR’s potential to generate abbreviated or shortcut names precludes the retrieval of this supplementary information. This discrepancy renders DDR not directly compatible with the retrieval-augmented autoformalizer framework in liu2025rethinking. To accommodate this, during our inference experiments, we set all required information fields to None, apart from the dependency item names generated by DDR. Despite this incompatibility, experimental results demonstrate that DDR still outperforms traditional retrieval methods. The primary contribution of DDR lies in offering an innovative and flexible solution for the field of autoformalization. It introduces a more direct and efficient approach to dependency retrieval and enables the flexible integration of various general-purpose LLMs as the core of RAG frameworks.
6 Conclusion
We identify three core limitations of dependency retrieval methods in current autoformalization research: low retrieval accuracy; a reliance on a pre-defined hyperparameter for the number of dependencies; and an inability to learn and improve from ever-growing datasets. To address these challenges, we propose a novel RAG paradigm, termed DDR (Direct Dependency Retrieval). DDR leverages the contextual understanding capabilities of LLMs to directly generate candidate dependencies for a given informal statement. This is coupled with an efficient verification mechanism based on suffix arrays to ensure the validity of the generated outputs. This approach obviates the need for a pre-set number of dependencies and is designed to continuously improve its performance by training on expanding public datasets. Our experimental results demonstrate that DDR significantly outperforms traditional retrieval methods in terms of both recall and precision on the dependency retrieval task, and it exhibits a corresponding performance advantage in the downstream autoformalization task.
Appendix A Examples Illustration
Informal statement example 1:
Prove that the total number of wooden blocks Thomas used to make the five stacks is $5$.
Formal statement example 1: theorem thm_P (a : $\mathbb{N}$) ($\text{h}_{0}$ : a > 0) ($\text{h}_{1}$ : a + (a + 1) + (a + 2) + (a + 3) + (a + 4) = 55) : a = 5 := by sorry Informal statement example 2:
Prove that the time it takes for Usain to run a whole lap around the school stadium is 42
Formal statement example 2: theorem thm_P : $\exists$ t, t > 0 $\wedge$ t = 42 := by sorry Informal statement example 3:
Prove that the number of ways $n$ tourists can distribute themselves among $n$ cinemas so that each cinema has exactly one tourist is $n!$.
Formal statement example 3: theorem thm_P (n : $\mathbb{N}$) : {f : Fin n $\rightarrow$ Fin n | f.Injective}.ncard = n! := by sorry Informal statement example 4:
Prove that the probability that James wins when he and his sister each spin a modified spinner with six congruent sectors numbered from 1 to 6, and James wins if the absolute difference of their numbers is 2 or less, is $\frac{2}{3}$.
Formal statement example 4: theorem thm_P : (Set.ncard {x : Fin 6 $\times$ Fin 6 | x.1 - x.2 $\in$ Finset.Icc (-2) 2}) / 36 = 2 / 3 := by sorry
Figure 6: Examples from the FineLeanCorpus with difficulty levels 0 and 1. These problems are phrased in a more natural-language-like style rather than formal mathematical statements, which makes them more suitable for ICL-based models.
Appendix B Experiment Details
For a fair comparison, we adopt an experimental setup identical to that of liu2025rethinking. We used the same prompt as described in Section A.7.1 of liu2025rethinking for evaluation on BEq. Similarly, the prompt for the ICL method is also consistent with that in Section A.7.3 of liu2025rethinking. DeepSeek-Prover-v2 (ren2025deepseek) is utilized as an auxiliary prover for BEq.
Our proposed DDR method is fine-tuned using the Swift (zhao2025swift) framework with the following hyperparameters:
  1. Max Sequence Length: 8192
  2. Batch size: 1
  3. Gradient Accumulation: 64
  4. Training Devices: 8
  5. Train Epochs: 6
  6. Optimizer: AdamW with learning rate $1\times 10^{-4}$, $\beta=(0.9,0.95)$, weight decay 0.1, maximal gradient norm 1.0, no warmup.
  7. Learning Rate Scheduler: Cosine.
  8. LoRA: rank 32, alpha 32, dropout 0.05.