### v4

# Improving Autoformalization Using Direct Dependency Retrieval

Abstract
Statement autoformalization, a crucial first step in formal verification, aims to transform informal descriptions of math problems into machine-verifiable formal representations but remains a significant challenge. The core difficulty lies in the fact that existing language models often suffer from a lack of contextual awareness, leading to hallucination of formal dependencies such as definitions and theorems. Current dependency retrieval approaches exhibit poor precision and recall, and lack the scalability to leverage ever-growing public datasets. To bridge this gap, we propose a novel retrieval-augmented framework based on Direct Dependency Retrieval (DDR). DDR directly generates candidate formal dependencies from natural-language mathematical descriptions and verifies their existence in the formal library via an efficient Suffix Array Check (SAC). Built on a SAC-constructed dependency retrieval dataset of over 500,000 samples, a high-precision DDR model is fine-tuned and shown to significantly outperform state-of-the-art methods in both retrieval precision and recall, leading to superior advantage in the autoformalization tasks. SAC also contributes in assessing formalization difficulty and enabling explicit quantification of the hallucination in In-Context Learning (ICL).
Improving Autoformalization Using Direct Dependency Retrieval
Shaoqi Wang   Lu Yu   Siwei Lou   Feng Yan Chunjie YangQing CuiJun Zhou Zhejiang University, Ant Group, Hong Kong Polytechnic University
1 Introduction
Theorem provers, such as Lean (moura2021lean), Coq (bertot2013interactive), and Isabelle (nipkow2002isabelle), are symbolic systems designed for formal verification whose soundness and completeness are proven. Grounded in specific foundational logics, these systems employ proprietary formal languages to encode mathematical statements, definitions, and proofs. The logical kernel within each system can then verify the validity of these formalized objects with complete rigor (nawaz2019survey). However, this process eschews the flexibility of natural language, demanding that users not only master the strict formal syntax but also possess a profound understanding of the semantic mapping between mathematical ideas and their formal representations. Consequently, the effective utilization of these theorem provers presents a formidable challenge to researchers (li2024survey).
Statement autoformalization refers to the process of automatically translating mathematical statements from natural language into a formal language comprehensible to theorem provers (liu2025rethinking). As a foundational step in the formal verification, the correctness of its output is paramount. It determines whether the proof target is precisely aligned with the intent of the original statement. Beyond this core function, statement autoformalization exhibits significant potential for broader applications and innovation. These applications include synthesizing training data for neural theorem provers (lin2025goedel), automated program verification (lin2024fvel), and curating and filtering informal reasoning datasets (zhang2025deeptheorem). , and evaluating and enhancing the informal reasoning capabilities of LLMs (e.g., through techniques like rejection sampling (zhou2024don)).
Current mainstream research in autoformalization primarily encompasses three direction. First, constructing larger and more challenging datasets for Supervised Fine-Tuning (SFT) (cao2025informal; peng2025criticlean; yu2025formalmath; miranda2025veribench; kimina_prover_2025; ying2024lean; zheng2021minif2f; azerbayev2023proofnet); second, encoding and indexing the libraries of theorem provers to support retrieval (liu2025rethinking; yang2023leandojo; wang2025aria); and third, employing methods like ICL to enable Large Language Models (LLMs) to directly translate informal statements into formal code (wu2022autoformalization; zhang2024consistent).
[…]
However, these approaches face two critical limitations. First, relying solely on generative models to directly synthesize formal code often overlooks the existing definitions and lemmas within the formal library (yang2023leandojo; zhang2024consistent). This ignorance is a primary reason for the model to hallucinate non-existent formal objects or generate syntactically incorrect code (wu2022autoformalization; liu2025rethinking). Results presented in Table 2 demonstrate that even with a two-stage ICL approach, where dependencies are retrieved prior to the main task, the problem of hallucination in model outputs remains significant. Second, the retrieval strategies for library dependencies in existing methods are rudimentary, yielding results with weak semantic relevance to the informal statement (azerbayev2023proofnet; yang2023leandojo). The libraries of mainstream theorem provers are vast and continually growing; for instance, Lean 4’s mathlib contains over 240,000 entries, including many with similar descriptions and usage. Simple retrieval methods struggle to achieve both high efficiency and precision. This challenge is confirmed by the experiments in liu2025rethinking as well as our own results presented in Table 3.
To address the aforementioned challenges, we propose a novel dependency retrieval method DDR, and the corresponding framework for statement autoformalization. In this context, dependency retrieval aims to identify relevant formal objects, such as definitions and theorems, from a theorem prover’s library based on an informal mathematical statement (liu2025rethinking). DDR leverages the contextual understanding of LLMs to directly generate potential formal dependencies. To empower the LLM with this capability, we have designed SAC for both data generation and dependency verification. Specifically, this process begins by constructing an efficient database from the theorem prover’s library using a suffix array. Subsequently, leveraging existing informal-formal statement pairs from public datasets, we extract candidate dependencies from the formal code and verify them against the database via the efficient binary search. This produces high-quality labels for fine-tuning the DDR model. The same verification mechanism is reused at inference time to validate the dependencies generated by the DDR model. Experimental results demonstrate that DDR significantly improves both recall and precision in dependency retrieval, leading to superior advantage in the autoformalization tasks.
Our main contributions are summarized as follows:
  1. 1.
    We propose DDR as a novel generation-plus-verification dependency retrieval paradigm in autoformalization tasks, and SAC as a flexible dependency verification method.
  2. 2.
    We provide a quantitative analysis of ICL for dependency retrieval, revealing its inherent and severe hallucination problems.
  3. 3.
    We conduct comprehensive experiments on statement autoformalization. The results validate that DDR can serve as a plug-and-play module to enhance the performance of existing autoformalizers and ICL methods.
  4. 4.
    The proposed SAC provides a potential sound mechanism for assessing formalization difficulty, overcomes the inherent subjectivity and stochasticity of LLM-based graders.
  5. 5.
    We have constructed and open-sourced a library dependency dataset with over 500,000 samples. The methodology for constructing this dataset is generalizable and can be applied to other formal corpora, fostering more diverse research in statement autoformalization.
2 Related Work
Formal mathematics aims to translate mathematical statements and proofs into machine-verifiable formal languages based on rigorous logical calculus. In recent years, with the advancement of deep learning, this field has become an active area of research, primarily encompassing two core branches: autoformalization (murphy2024autoformalizing; liu2025rethinking; poiroux2024improving; azerbayev2023proofnet; wang2020exploration), and automated theorem proving (lin2025goedel; kimina_prover_2025; chen2025seed; xin2024deepseek; xin2024deepseekv15; ren2025deepseek; liu2025proofaug).
Figure 2: Schematic diagram of SAC process. All library items are concatenated into a single string with delimiters, followed by suffix array construction. Each pending item is then queried via binary search to determine its match status (exact, partial, or none). This mechanism supports both dataset construction and DDR output verification.
Automated theorem proving is dedicated to automatically generating proof scripts for given formal goals. The central challenge lies in designing efficient proof search strategies (liu2025proofaug). Current cutting-edge research in this direction focuses on combining LLMs with reinforcement learning to emulate the heuristic reasoning processes of human mathematicians in exploring the proof space (chen2025seed), with the future prospect of assisting in the discovery of new theorems. Prominent examples of such systems include Seed-Prover (chen2025seed) and DeepSeek-Prover (xin2024deepseek; xin2024deepseekv15; ren2025deepseek). In contrast, autoformalization targets the automatic translation of informal mathematical content, typically expressed in natural language, into a formal language. Research in this area concentrates on establishing reliable evaluation methodologies to assess the semantic fidelity of the translation (liu2025rethinking; poiroux2024improving) and on developing higher-performance autoformalization models (liu2025rethinking; poiroux2024improving). This paper focuses on the task of statement autoformalization. We propose a method to mitigate model hallucination and enhance model performance by offering better dependency retrieval results in the RAG framework.
Progress in autoformalization research has led to the creation of large-scale informal-to-formal parallel corpora (mahdavi2025leveraging). These corpora, such as FormalMath (yu2025formalmath), CriticBench (peng2025criticlean), and FineLeanCorpus (peng2025criticlean), have expanded in scale from a few hundred samples to tens of thousands and are primarily generated through rule-based methods and by leveraging the ICL capabilities of LLMs (azerbayev2023proofnet; gao2024omni). However, current autoformalization research has not effectively utilized this data to discover the intricate connection between informal mathematical statements and their required library dependencies. To address this gap, we propose DDR, a method designed to enhance a model’s ability to infer and utilize library dependencies during statement formalization.
Research on evaluating the fidelity of autoformalization has explored various methods, such as manual expert review (wang2018first; peng2025criticlean), LLM review (yang2023leandojo; sidhu2024evaluation), auxiliary metrics like Perplexity (wang2018first), BLEU (papineni2002bleu; wang2018first), and type checking (poiroux2024improving) within a theorem prover. However, these approaches have inherent limitations in capturing the deep semantic equivalence between informal and formal mathematical statements. They may overlook semantic inconsistencies between the two modalities or produce false negatives by incorrectly rejecting semantically valid formalizations (liu2025rethinking). To overcome this, our work adopts Bidirectional Extender Definitional Equivalence (BEq) as the primary evaluation metric (liu2025rethinking). This choice is motivated by BEq’s high alignment with human expert judgment in assessing semantic equivalence. The core principle of BEq is to define two formal statements as equivalent if they are mutually transformable through a predefined, finite, and verifiable set of transformation rules (aczel2013homotopy).
3 Methodology
3.1 Direct Dependency Retrieval
To prevent hallucination during generation, existing dependency retrieval methods select items from a pre-existing library to form their output (yang2023leandojo; liu2025rethinking), as illustrated in Figure 1 (a). However, this selected-based paradigm suffers from two primary limitations. First, training an encoder to produce effective embeddings for every formal object in the library to maximize retrieval performance is a challenge in itself (xiong2020approximate; gao2023retrieval). Second, the method is sensitive to superficial variations in the input: semantically equivalent mathematical statements may yield different results in maximum cosine similarity retrieval due to minor differences in their syntactic structure or notational representation (asl2024robustsentembed). Our empirical results in Table 3 confirm that such embedding-based retrieval methods exhibit low recall and precision. Inaccurate retrieval results don’t contribute to the language model in autoformalization tasks. Figure 1 (a) illustrates this failure mode: in the absence of a usable dependency from the retrieved Mathlib candidates, the model attempts to construct a local function functionally equivalent to factorization, leading to an erroneous output. Furthermore, this approach lacks scalability, as it cannot effectively leverage the growing volume of available datasets to continuously enhance model capabilities.
Difficulty Num Depend Rate Depend Length 0 835 0.188 023 952 095 808 380.188\,023\,952\,095\,808\,38 0.267 065 868 263 473 060.267\,065\,868\,263\,473\,06 1 120040 0.555 081 639 453 515 50.555\,081\,639\,453\,515\,5 1.061 096 301 232 922 41.061\,096\,301\,232\,922\,4 2 113816 0.783 756 238 138 750 30.783\,756\,238\,138\,750\,3 1.862 936 669 712 518 51.862\,936\,669\,712\,518\,5 3 49841 0.807 347 364 619 490.807\,347\,364\,619\,49 2.079 312 212 836 8212.079\,312\,212\,836\,821 4 69882 0.795 240 548 352 937 80.795\,240\,548\,352\,937\,8 1.957 700 123 064 594 51.957\,700\,123\,064\,594\,5 5 67547 0.523 383 718 003 760 40.523\,383\,718\,003\,760\,4 1.105 689 371 844 789 51.105\,689\,371\,844\,789\,5 6 41312 0.810 878 195 197 521 40.810\,878\,195\,197\,521\,4 2.229 908 985 282 726 52.229\,908\,985\,282\,726\,5 7 36675 0.753 946 830 265 848 60.753\,946\,830\,265\,848\,6 2.108 166 325 835 037 42.108\,166\,325\,835\,037\,4 8 8682 0.819 511 633 264 224 80.819\,511\,633\,264\,224\,8 2.393 112 186 132 227 82.393\,112\,186\,132\,227\,8 9 708 0.824 858 757 062 146 80.824\,858\,757\,062\,146\,8 2.611 581 920 903 9552.611\,581\,920\,903\,955
Table 1: Dataset statistics of FineLeanCorpus. Depend Rate denotes the percentage of samples that have explicit mathlib dependencies within each difficulty level. Depend Length denotes the mean number of explicit mathlib dependency items for samples within each difficulty level. Higher Depend Rate and Depend Length values typically indicate more challenging tasks.
Method    Diff01    Diff23    Diff45    Diff67    Diff89 $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ $_{\text{m}} Hall$$_{\text{std}}$ Claude-3.5-Sonnet 0.349 842 767 296 226 50.349\,842\,767\,296\,226\,5 0.321 705 896 805 355 60.321\,705\,896\,805\,355\,6 0.300 598 443 2230.300\,598\,443\,223 0.257 305 310 183 968 60.257\,305\,310\,183\,968\,6 0.277 148 809 5230.277\,148\,809\,523 0.258 708 434 626 845 840.258\,708\,434\,626\,845\,84 0.284 793 650 793 500 050.284\,793\,650\,793\,500\,05 0.220 764 928 185 668 670.220\,764\,928\,185\,668\,67 0.276 160 173 1630.276\,160\,173\,163 0.221 181 722 747 262 330.221\,181\,722\,747\,262\,33 DeepSeek-R1 0.234 589 265 257 731 940.234\,589\,265\,257\,731\,94 0.282 361 595 375 512 250.282\,361\,595\,375\,512\,25 0.330 162 841 356 923 10.330\,162\,841\,356\,923\,1 0.313 336 322 060 300 80.313\,336\,322\,060\,300\,8 0.265 354 183 465 656 60.265\,354\,183\,465\,656\,6 0.289 163 096 881 404 20.289\,163\,096\,881\,404\,2 0.306 544 110 086 50.306\,544\,110\,086\,5 0.297 199 460 504 328 20.297\,199\,460\,504\,328\,2 0.302 954 725 8290.302\,954\,725\,829 0.288 580 908 324 351 80.288\,580\,908\,324\,351\,8 GPT-4o 0.290 530 303 031 818 20.290\,530\,303\,031\,818\,2 0.279 890 001 371 935 650.279\,890\,001\,371\,935\,65 0.298 034 188 031 282 070.298\,034\,188\,031\,282\,07 0.274 865 951 589 759 10.274\,865\,951\,589\,759\,1 0.287 544 867 191 959 80.287\,544\,867\,191\,959\,8 0.265 736 723 207 650 960.265\,736\,723\,207\,650\,96 0.290 236 898 778 392 040.290\,236\,898\,778\,392\,04 0.278 007 943 706 211 30.278\,007\,943\,706\,211\,3 0.300 733 890 125 757 630.300\,733\,890\,125\,757\,63 0.258 576 099 806 656 40.258\,576\,099\,806\,656\,4 Ling-flash-2.0 0.326 612 785 787 1560.326\,612\,785\,787\,156 0.316 176 237 531 755 40.316\,176\,237\,531\,755\,4 0.385 175 652 200 500 040.385\,175\,652\,200\,500\,04 0.314 393 792 807 857 730.314\,393\,792\,807\,857\,73 0.342 384 920 637 000 10.342\,384\,920\,637\,000\,1 0.298 188 535 254 873 270.298\,188\,535\,254\,873\,27 0.325 408 019 2680.325\,408\,019\,268 0.276 772 886 385 413 10.276\,772\,886\,385\,413\,1 0.362 064 913 029 000 070.362\,064\,913\,029\,000\,07 0.276 638 773 854 510 90.276\,638\,773\,854\,510\,9 Qwen-Max 0.229 807 692 306 730 760.229\,807\,692\,306\,730\,76 0.302 868 995 190 336 670.302\,868\,995\,190\,336\,67 0.274 352 489 913 265 330.274\,352\,489\,913\,265\,33 0.305 769 901 832 063 040.305\,769\,901\,832\,063\,04 0.209 369 488 533 333 340.209\,369\,488\,533\,333\,34 0.241 521 686 559 3190.241\,521\,686\,559\,319 0.241 563 880 005 025 130.241\,563\,880\,005\,025\,13 0.264 552 600 591 864 860.264\,552\,600\,591\,864\,86 0.247 436 120 550 753 790.247\,436\,120\,550\,753\,79 0.238 268 638 411 097 50.238\,268\,638\,411\,097\,5 R@5 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 R@10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 DDR $0.016$ 0.083 891 822 234 576 820.083\,891\,822\,234\,576\,82 0.006 027 253 668 553 460.006\,027\,253\,668\,553\,46 0.036 863 586 454 198 930.036\,863\,586\,454\,198\,93 0.004 540 420 820 155 0380.004\,540\,420\,820\,155\,038 0.026 243 591 065 426 320.026\,243\,591\,065\,426\,32 0.003 062 678 062 820 512 80.003\,062\,678\,062\,820\,512\,8 0.022 527 428 053 560 6230.022\,527\,428\,053\,560\,623 0.013 735 240 031 097 560.013\,735\,240\,031\,097\,56 0.087 415 728 148 786 620.087\,415\,728\,148\,786\,62
[…]
To address the challenge of LLMs hallucinating non-existent or irrelevant library dependencies during autoformalization, we propose a generation-plus-verification pipeline, with DDR as its core component. DDR is a SFT model to operate in an end-to-end fashion, directly extracting potential library dependencies from an input mathematical statement. The SFT based paradigm enables it to effectively absorb and leverage continuously growing training data, thereby demonstrating excellent scalability. In contrast to conventional two-stage retrieval methods that require additional encoder model (liu2025rethinking), the design of DDR streamlines the architecture, leading to a simpler, more efficient, and better-performing approach. Accurate dependency retrieval reduces the complexity of the subsequent formalization task, thereby enhancing the overall performance of the autoformalizer. As illustrated in Figure 1 (b), providing the model with precise dependencies enables it to more reliably generate the correct formal output, thus increasing the overall success rate.
The effective application of the DDR method entails two primary challenges: first, constructing a high-quality dataset for model fine-tuning, and second, implementing an output filtering mechanism to validate the generated dependencies. Such challenges can be addressed by the following SAC method.
3.2 Suffix Array Check
To accurately and efficiently verify the existence of dependency objects generated by LLMs within a formal library, we introduce a checking method based on suffix arrays (liu2024infini). The core idea of this approach is to first concatenate the identifiers of all defined objects in the library, such as pre-proven theorems, into a single long string using a specific delimiter. Subsequently, a suffix array is constructed from this concatenated long string in linear complexity (lee2022deduplicating; karkkainen2006linear). By definition, a suffix array is an array of integers that stores the starting indices of all suffixes of the original string, sorted in lexicographical order (karkkainen2006linear). Figure 2 illustrates the SAC process with a simplified example, assuming the library contains only two objects and there are three pending items. The sorted nature of the suffix array facilitates efficient matching of these pending items via binary search. Specifically, for each pending item, we can locate the range of all suffixes that have the item as a prefix within the suffix array. This process enables the rapid retrieval of all library objects corresponding to the pending items, thereby completing their existence verification.
SAC is an efficient method for verification, capable of meeting the real-time requirements of large-scale requests (liu2024infini). For a formal library containing $N$ objects with an average length of $d$, and $M$ pending items with an average length of $s$, conventional string matching methods typically involve brute-force comparison, resulting in a time complexity of $\mathcal{O}\left((s+d)MN\right)$. In contrast, the suffix array auxiliary check operates on a suffix array constructed over the entire library. For each pending item, it utilizes binary search on the suffix array to locate matches. The overall time complexity of this approach is $\mathcal{O}(sM\log N)$. This significant reduction in computational complexity enables real-time responsiveness (liu2024infini).
3.3 Constructing High-quality Training Dataset
Fine-tuning a precise and efficient DDR model necessitates a dataset of pairs, each comprising a mathematical statement and its corresponding library dependencies. However, no such public dataset is currently available, and its construction is primarily challenged by the difficulty of accurately extracting and efficiently verifying library dependencies from existing formal statements. To address this, we employ the above SAC method to process a large-scale autoformalization corpus to generate the required training data. We select FineLeanCorpus (peng2025criticlean) as the source corpus, which contains over 500,000 samples spanning a wide range of mathematical domains and difficulty levels. After partitioning the corpus into training and test sets based on problem difficulty, we leverage the efficiency of SAC to process the entire training partition in under two hours. This process yield a high-quality dataset suitable for fine-tuning the DDR model. This resulting dataset has been made open-access at https://huggingface.co/datasets/Palca/FineLeanCorpus_DDR. We anticipate that this contribution will foster further research on the field of formal mathematical proving.
4 Direct Dependency Retrieval-augmented Autoformalization
Method    Diff01    Diff23    Diff45    Diff67    Diff89 Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Claude-3.5-Sonnet 0.490 249 999 999 999 960.490\,249\,999\,999\,999\,96 $0.5235$ 0.277 160 714 285 714 260.277\,160\,714\,285\,714\,26 0.384 761 904 761 904 750.384\,761\,904\,761\,904\,75 $0.214$ 0.340 333 333 333 333 30.340\,333\,333\,333\,333\,3 0.279 833 333 333 333 30.279\,833\,333\,333\,333\,3 0.434 710 317 460 317 40.434\,710\,317\,460\,317\,4 0.272 853 354 978 354 950.272\,853\,354\,978\,354\,95 0.333 990 079 365 079 40.333\,990\,079\,365\,079\,4 DeepSeek-R1 0.531 833 333 333 333 30.531\,833\,333\,333\,333\,3 0.557 833 333 333 333 30.557\,833\,333\,333\,333\,3 0.240 378 663 003 6630.240\,378\,663\,003\,663 0.353 684 523 809 523 750.353\,684\,523\,809\,523\,75 0.193 174 603 174 603 170.193\,174\,603\,174\,603\,17 0.323 297 619 047 6190.323\,297\,619\,047\,619 0.230 887 371 615 312 780.230\,887\,371\,615\,312\,78 0.368 638 888 888 888 850.368\,638\,888\,888\,888\,85 0.262 246 031 746 031 70.262\,246\,031\,746\,031\,7 0.371 527 777 777 777 80.371\,527\,777\,777\,777\,8 GPT-4o 0.538 750 000 000 000 10.538\,750\,000\,000\,000\,1 0.548 083 333 333 333 40.548\,083\,333\,333\,333\,4 0.282 880 952 380 952 40.282\,880\,952\,380\,952\,4 0.317 470 238 095 238 10.317\,470\,238\,095\,238\,1 0.203 500 000 000 000 010.203\,500\,000\,000\,000\,01 0.247 499 999 999 999 970.247\,499\,999\,999\,999\,97 0.262 416 666 666 666 70.262\,416\,666\,666\,666\,7 0.318 698 412 698 412 70.318\,698\,412\,698\,412\,7 0.274 305 555 555 555 50.274\,305\,555\,555\,555\,5 0.289 980 158 730 158 750.289\,980\,158\,730\,158\,75 Ling-flash-2.0 0.449 250.449\,25 0.469 333 333 333 333 40.469\,333\,333\,333\,333\,4 0.173 851 190 476 190 50.173\,851\,190\,476\,190\,5 0.236 178 571 428 571 430.236\,178\,571\,428\,571\,43 0.147 541 666 666 666 650.147\,541\,666\,666\,666\,65 0.219 499 999 999 999 950.219\,499\,999\,999\,999\,95 0.214 780 122 655 122 670.214\,780\,122\,655\,122\,67 0.319 948 412 698 412 70.319\,948\,412\,698\,412\,7 0.181 663 295 038 2950.181\,663\,295\,038\,295 0.247 426 587 301 587 310.247\,426\,587\,301\,587\,31 Qwen-Max 0.479 464 285 714 285 70.479\,464\,285\,714\,285\,7 0.511 250.511\,25 0.263 158 730 158 730 150.263\,158\,730\,158\,730\,15 0.293 380 952 380 952 40.293\,380\,952\,380\,952\,4 0.227 833 333 333 333 30.227\,833\,333\,333\,333\,3 0.276 583 333 333 333 350.276\,583\,333\,333\,333\,35 0.262 250.262\,25 0.331 380 952 380 952 40.331\,380\,952\,380\,952\,4 0.266 392 857 142 857 10.266\,392\,857\,142\,857\,1 0.323 132 936 507 936 450.323\,132\,936\,507\,936\,45 R@5 0.022 000 000 000 000 0020.022\,000\,000\,000\,000\,002 0.069 166 666 666 666 650.069\,166\,666\,666\,666\,65 $0.073$ 0.150 970 238 095 238 070.150\,970\,238\,095\,238\,07 0.048 999 999 999 999 9950.048\,999\,999\,999\,999\,995 0.119 250 000 000 000 010.119\,250\,000\,000\,000\,01 $0.079$ 0.171 166 666 666 666 660.171\,166\,666\,666\,666\,66 $0.085$ 0.143 722 222 222 222 220.143\,722\,222\,222\,222\,22 R@10 0.016 500 000 000 000 0040.016\,500\,000\,000\,000\,004 0.100 416 666 666 666 650.100\,416\,666\,666\,666\,65 0.044 499 999 999 999 990.044\,499\,999\,999\,999\,99 0.192 136 904 761 904 760.192\,136\,904\,761\,904\,76 $0.032$ 0.150 583 333 333 333 320.150\,583\,333\,333\,333\,32 0.059 500 000 000 000 0040.059\,500\,000\,000\,000\,004 0.265 476 190 476 190 530.265\,476\,190\,476\,190\,53 $0.061$ 0.206 480 158 730 158 70.206\,480\,158\,730\,158\,7 DDR 0.84 0.82 0.90 0.87 0.91 0.92 0.91 0.88 0.83 0.83
Table 3: Comparison of the retrieval dependency results. Results shown are filtered (i.e., without hallucinations).
Figure 3: LLM consistency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
This section presents a comprehensive evaluation of the proposed method. Our evaluation is two-fold: first, we directly assess the performance of DDR on the dependency retrieval task. Second, we measure the improvements it brings by evaluating the pass@8 success rate on the downstream autoformalization task. Furthermore, the experiments provide additional insights, such as a discussion on hallucinations in ICL methods and an analysis of the difficulty of autoformalization datasets.
4.1 Dataset Overview and Difficulty Analysis
The difficulty attribute is a key metric in FineLeanCorpus designed to quantify the challenge of formalization, rated on an 11-point scale from 0 to 10. In our experimental setup, we consolidated samples with a difficulty of 10 into the difficulty 9 category. The statistical results of Mathlib dependency on datasets are summarized in Table 1.
We randomly sampled 100 statements from each difficulty category and then combined the samples from every two adjacent difficulty categories as the test set. This process resulted in five test sets, named Diff01, Diff23, Diff45, Diff67, and Diff89, where Diffij denotes the combined set of samples from difficulty levels i and j, with each test set containing 200 statements. All remaining samples in the dataset were used to fine-tune the DDR model.
The original difficulty ratings for the FineLeanCorpus (peng2025criticlean) datasets, assigned by an LLM grader, exhibit a notable discrepancy with our experimental results in Table 4 and Table 5. To investigate this inconsistency, we propose assessing the actual formalization difficulty using two alternative metrics: Depend Rate and Depend Length. Based on the data in Table 1, we can calculate the weighted averages of Depend Rate and Depend Length for the Diff23, Diff45, Diff67, and Diff89 subsets. The results are as follows: Diff45 has the lowest values (0.662, 1.539), the values for Diff23 (0.791, 1.929) and Diff67 (0.784, 2.173) are higher, and Diff89 has the highest values (0.820, 2.410). These metrics reveal a difficulty ranking of Diff89 > Diff67 > Diff23 > Diff45. This ranking aligns closely with our experimental observations, suggesting that Depend Rate and Depend Length, derived via our DDR and SAC, can serve as a potential indicator of formalization difficulty along with LLM-based graders. The Diff01 subset is excluded from this analysis because its imprecise informal descriptions introduce extraneous formalization challenges, making its difficulty hard to assess with these metrics. Detailed discussion are presented in appendix B
Figure 4: LLM efficiency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq MMA $0.47$ $0.095$ $0.455$ $0.11$ $0.6$ $0.19$ $0.5$ $0.1$ $0.385$ $0.065$ RA $0.67$ $0.115$ $0.730$ $0.170$ $0.765$ $0.295$ $0.735$ $0.185$ $0.61$ $0.105$ RA+R@5 $0.86$ $0.185$ $0.795$ $0.18$ $0.805$ $0.145$ $0.825$ $0.115$ $0.765$ $0.135$ RA+R@10 0.880 $0.205$ $0.86$ $0.2$ 0.910 $0.255$ $0.825$ $0.16$ $0.765$ 0.150 RA+DDR $0.865$ 0.235 0.900 0.215 0.910 0.305 0.865 0.225 0.770 0.150
Table 4: Comparison of the autoformalization results using SFT methods. MMA denotes the LLM fine-tuned on MMA’s Lean subset. RA refers to the pure autoformalization method from liu2025rethinking, while R@5 and R@10 denote its retrieval-augmented variants using the top-5 and top-10 retrieved results, respectively. RA+DDR denotes the new retrieval-augmented variant incorporating the proposed DDR.
4.2 Setup
Due to space constraints, we defer the detailed description of our experimental setup to the appendix C. Our evaluation utilizes two primary metrics: Type Checking (TypeC) and Bidirectional Extended Definitional Equality (BEq). We compare our method against a comprehensive set of baseline methods, including MMA (jiang2023multilingual), DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (anthropic_claude35_sonnet_2024), GPT-4o (achiam2023gpt), Qwen-Max (qwen2025qwen25technicalreport), as well as RA+R@5 and RA+R@10 (liu2025rethinking).
4.3 Hallucination Study
Prior to the introduction of DDR, a sound mechanism for assessing dependency hallucinations in the outputs of LLMs on autoformalization tasks was lacking. The proposed DDR framework, through the integrated SAC, provides a practical and efficient solution to this problem. This method determines whether each generated dependency exists within the formal library, thereby enabling the quantification of hallucination severity. Define the hallucination rate (Hall) as the proportion of hallucinatory items in a given set of dependencies. For instance, if a model outputs (Real.div, Real.max, Nat.succ) and only Nat.succ is authentic, the Hall is 2/3. $\textbf{Hall}_{\textbf{m}}$ and $\textbf{Hall}_{\textbf{std}}$ denote the mean and standard deviation of this rate across the test set, respectively. The hallucination rates for various mainstream dependency retrieval methods are presented in Table 2.
The analysis reveals that R@5 and R@10, as select-based methods, are inherently free from hallucination by design, since their outputs are drawn directly from the formal library. In contrast, all generative retrieval methods are susceptible to hallucination. Our proposed DDR method exhibits outstanding performance, with a mean hallucination rate ($\textbf{Hall}_{\textbf{m}}$) below 0.02 and a standard deviation ($\textbf{Hall}_{\textbf{std}}$) below 0.09, indicating nearly non-existent hallucination. In stark contrast, other ICL-based methods suffer from severe hallucination, with both the mean and standard deviation of their Hall values hovering around 0.30. This implies that, on average, approximately 30% of the dependencies generated by these methods are hallucinatory.
Method RAG    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq Claude-3.5-Sonnet N/A $0.83$ $0.415$ $0.74$ $0.33$ 0.775 $0.385$ 0.680 $0.205$ $0.53$ $0.1$ Claude-3.5-Sonnet ICL $0.74$ $0.37$ $0.515$ $0.245$ $0.64$ $0.33$ $0.455$ $0.145$ $0.385$ $0.11$ Claude-3.5-Sonnet DDR 0.905 0.460 0.770 0.335 $0.765$ 0.425 $0.625$ 0.220 0.630 0.130 DeepSeek-R1 N/A $0.91$ $0.225$ $0.745$ $0.205$ $0.82$ $0.36$ $0.78$ $0.23$ $0.66$ $0.13$ DeepSeek-R1 ICL $0.88$ $0.34$ $0.755$ $0.315$ $0.825$ $0.44$ $0.76$ $0.27$ $0.57$ $0.135$ DeepSeek-R1 DDR 0.940 0.410 0.875 0.385 0.940 0.515 0.875 0.345 0.695 0.200 GPT-4o N/A $0.905$ $0.44$ $0.76$ $0.295$ $0.8$ $0.4$ $0.745$ $0.24$ $0.575$ $0.135$ GPT-4o ICL $0.88$ 0.460 $0.675$ $0.255$ $0.73$ $0.345$ $0.625$ $0.21$ $0.405$ $0.12$ GPT-4o DDR 0.940 0.460 0.775 0.320 0.850 0.455 0.780 0.275 0.630 0.150 Ling-flash-2.0 N/A 0.850 0.405 0.610 $0.21$ 0.665 $0.3$ $0.51$ $0.19$ $0.38$ $0.085$ Ling-flash-2.0 ICL $0.775$ $0.355$ $0.42$ $0.17$ $0.59$ $0.28$ $0.395$ $0.17$ $0.295$ $0.085$ Ling-flash-2.0 DDR $0.775$ $0.4$ $0.595$ 0.225 $0.63$ 0.360 0.535 0.205 0.440 0.115 Qwen-Max N/A $0.795$ 0.375 $0.605$ $0.22$ 0.665 0.360 $0.62$ $0.205$ $0.44$ $0.105$ Qwen-Max ICL $0.73$ $0.315$ $0.445$ $0.16$ $0.55$ $0.26$ $0.445$ $0.165$ $0.35$ $0.115$ Qwen-Max DDR 0.830 0.375 0.660 0.250 $0.56$ $0.285$ 0.695 0.255 0.545 0.120
Table 5: Comparison of the autoformalization results using ICL methods. RAG column indicates the dependency retrieval strategy used in autoformalization: N/A denotes no retrieval, ICL denotes prompt-based retrieval, and DDR refers to the proposed method.
4.4 Retrieval Results
Although R@5 and R@10 are hallucination-free, their retrieval performance is questionable. To isolate retrieval quality from the issue of hallucination and enable a fair comparison of the methods’ true capabilities, we first filtered out all hallucinatory items from each method’s output and then recalculated their precision and recall. The performance comparison after filtering is shown in Table 3.
After filtering, the results show substantial performance disparities among the models. Despite having no hallucinations, R@5 and R@10 yield the lowest precision (0.02-0.10) and recall (0.10-0.20), indicating their retrieval capability is severely limited. The filtered prompt-based methods achieve moderate performance, with precision scores ranging from 0.2 to 0.5 and recall scores from 0.3 to 0.5. Our proposed DDR method demonstrates the most superior performance, achieving a high precision of 0.8-0.9 and a recall of 0.8-0.9, significantly outperforming all other baseline methods.
4.5 Autoformalization Results
The efficacy of DDR is further substantiated by its performance improvements in the downstream autoformalization task. As shown in Table 4 and Table 5, integrating DDR yields significant performance gains under both SFT and ICL frameworks. We specifically evaluated the combination of DDR with the RA model, an advanced model fine-tuned for autoformalization in liu2025rethinking. A format mismatch exists, as the RA model requires a comprehensive input format (full dependency name, informal description, and code), whereas DDR outputs only abbreviated dependency names. Consequently, the R@5 and R@10 retrievers from liu2025rethinking, designed for RA, are more format-compatible. Despite this mismatch, the combination of RA and DDR consistently outperforms all baseline methods. This success is primarily attributed to DDR’s superior retrieval precision and recall, highlighting its excellent generalization and robustness. In contrast, the performance of R@5 and R@10 reflects a trade-off between precision and recall, consistent with their respective retrieval characteristics.
To evaluate DDR’s potential as a plug-and-play component, we integrated it with ICL methods. Table 5 compares three experimental setups: (1) direct autoformalization, (2) two-stage autoformalization using a prompt-based dependency retrieval, and (3) two-stage autoformalization using a DDR-based dependency retrieval. The results indicate that the DDR-based RAG is the superior strategy in most test scenarios. Experiments on Deepseek-R1, illustrated in Figure 3, further reveal DDR’s stability advantage: the DDR-integrated method not only achieves a higher total pass@8 success count but also exhibits a significantly better distribution of successful attempts across 8 independent trials compared to the baseline. Furthermore, Figure 4 shows that DDR can complete more tasks with fewer attempts.
5 Conclusion
We propose DDR, a method that leverages the contextual understanding of LLMs to directly generate candidate dependencies for a given informal statement in autoformalization tasks. This is complemented by SAC, which utilize suffix array for efficient dependency verification. Experimental results validate the effectiveness of our approach and demonstrate its potential in ICL hallucinations analysis and in assessing the difficulty of autoformalization datasets.
6 Limitations
A key limitation of DDR is its inability to precisely locate the generated items within the library. Specifically, whereas the full retrieval results in methods like liu2025rethinking typically include complete metadata such as the item’s name, code usage, and documentation, DDR’s potential to generate abbreviated or shortcut names precludes the retrieval of this supplementary information. This discrepancy renders DDR not directly compatible with the retrieval-augmented autoformalizer framework in liu2025rethinking.
DDR also exhibits limitations in handling certain special cases, which necessitate dedicated mechanisms. For instance, some core constructs (such as MOD), despite being crucial and frequently used in formalization, may not exist as standard library items
Formal mathematics libraries are continuously evolving, with dependency items being added, removed, or renamed across different versions. Such changes require corresponding maintenance of our system to ensure its continued functionality and effectiveness.
Appendix A Additional Results
A.1 Full Figures
Figure 5: LLM consistency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
Figure 6: LLM efficiency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
While Figure 3 and Figure 4 primarily presented the experimental results on the Deepseek-R1 model, this section provides a more comprehensive evaluation by integrating the DDR method with various other ICL methods in Figure 5 and Figure 6. The experimental data demonstrates that incorporating DDR brings significant performance improvements to these baseline methods in the vast majority of cases.
A.2 Out of Domain Study
To evaluate the generalization ability of DDR, we conduct autoformalization experiments on the Omni-math dataset (gao2024omni), adopting the experimental setup from lin2025goedel. We compare DDR against the baseline methods Goedel-Formalizer-V2-8B (lin2025goedel) and Kimina-Autoformalizer-7B (kimina_prover_2025). To evaluate the correctness of the generated formalizations, we employ the method proposed in lin2025goedel, which combines Lean4 TypeC with semantic judgments from LLMs. The pass@8 results are presented in Table 6. The results demonstrate that DDR enhances the performance of existing autoformalizers, even those not specifically designed for dependency retrieval.
Name RAG Pass Failed Goedel-Formalizer-V2-8B N/A 180 120 Goedel-Formalizer-V2-8B DDR 183 117 Kimina-Autoformalizer-7B N/A 105 195 Kimina-Autoformalizer-7B DDR 128 172
Table 6: Comparison of different formalizers on 300 Omni-math problems.
Appendix B Examples Illustration of Diff01
A notable observation is that samples categorized as difficulty 0 and difficulty 1 present a counter-intuitively high formalization challenge. This is primarily because the natural language statements in this category often do not represent abstract descriptions of the problem’s mathematical essence, as illustrated by examples in Figure 7. Our subsequent experimental results corroborate this finding: for Diff01 dataset, the performance of autoformalization across all methods are worse than Diff23 and Diff45, yet still better than that of Diff89.
Figure 7 presents several examples from the Diff01 dataset. These problems are primarily formulated in a natural language style rather than as rigorous formal mathematical statements. Such presentation brings additional challenge in autoformalization, and renders difficulty metrics that rely on Depend Rate and Depend Length inapplicable, thereby impeding an accurate quantification of the autoformalization task difficulty on this dataset. Furthermore, because these problem descriptions lack an abstract representation of their mathematical essence, all methods evaluated in Table 4 and Table 5 exhibit a significant degradation in performance on the Diff01 dataset.
Informal statement example 1:
Prove that the total number of wooden blocks Thomas used to make the five stacks is $5$.
Formal statement example 1: theorem thm_P (a : $\mathbb{N}$) ($\text{h}_{0}$ : a > 0) ($\text{h}_{1}$ : a + (a + 1) + (a + 2) + (a + 3) + (a + 4) = 55) : a = 5 := by sorry Informal statement example 2:
Prove that the time it takes for Usain to run a whole lap around the school stadium is 42
Formal statement example 2: theorem thm_P : $\exists$ t, t > 0 $\wedge$ t = 42 := by sorry Informal statement example 3:
Prove that the number of ways $n$ tourists can distribute themselves among $n$ cinemas so that each cinema has exactly one tourist is $n!$.
Formal statement example 3: theorem thm_P (n : $\mathbb{N}$) : {f : Fin n $\rightarrow$ Fin n | f.Injective}.ncard = n! := by sorry Informal statement example 4:
Prove that the probability that James wins when he and his sister each spin a modified spinner with six congruent sectors numbered from 1 to 6, and James wins if the absolute difference of their numbers is 2 or less, is $\frac{2}{3}$.
Formal statement example 4: theorem thm_P : (Set.ncard {x : Fin 6 $\times$ Fin 6 | x.1 - x.2 $\in$ Finset.Icc (-2) 2}) / 36 = 2 / 3 := by sorry
Figure 7: Examples from the FineLeanCorpus with difficulty levels 0 and 1.
Appendix C Experiment Settings
C.1 SAC Details
We employ precision and recall as quantitative metrics to evaluate the performance of the dependency retrieval results. A key challenge in this evaluation arises from the discrepancy between the outputs of DDR, which directly generate dependency short names, and the ground-truth labels, which are Mathlib object identifiers extracted from source code. This leads to name ambiguity, as source code often contains unqualified or partially qualified names. To accurately identify true positives for our evaluation, we introduce a suffix-based matching criterion. Specifically, we represent each identifier as a sequence of dot-separated components. A retrieved identifier is considered a match for a ground-truth label if and only if two conditions are met: (1) the string representation of one identifier is a substring of the other; and (2) when both identifiers are split by the dot character (.) into lists of their components, one list is a sublist of the other.
C.2 Metrics
To evaluate the semantic equivalence between the formal statements predicted by the model and the ground-truth formal statements, we employ two criteria: Type Checking (TypeC) and Bidirectional Extended Definitional Equality (BEq (liu2025rethinking). TypeC ensures the syntactic correctness of expressions, whereas BEq is utilized to determine their semantic equivalence. Building upon these criteria, we define the TypeC@8 and BEq@8 as our final evaluation metrics shown in Table 4 and Table 5. These metrics measure the percentage of problems for which at least one of the top-8 predicted statements is deemed equivalent to the ground-truth statement under the corresponding criterion.
C.3 Hallucination Details
[…]
C.4 Baseline Details
To comprehensively evaluate our proposed method, we established several baseline models and conducted ablation studies. For the ablation studies, we employed the RAutoformalizer (RA) from liu2025rethinking as a pure formalizer baseline, and RA+R@k as an autoformalizer that incorporates the top-k retrieved dependencies using the method in liu2025rethinking. The RA was fine-tuned based on InternLM2-Math-Base-7B (ying2024internlm). Our baselines cover multiple aspects: first, we evaluated a range of state-of-the-art general LLMs via ICL, including DeepSeek-R1 (guo2025deepseek), Ling-flash-2.0 (ai2025ming), Claude-3.5-Sonnet (claude-3-5-sonnet-20241022) (anthropic_claude35_sonnet_2024), GPT-4o (gpt-4o-2024-08-06) (hurst2024gpt), and Qwen-Max (qwen-max-2024-09-19) (team2024qwen2); second, we included an InternLM2-Math-Base-7B (ying2024internlm) model fine-tuned on the MMA Lean subset (jiang2023multilingual). For our proposed DDR method, we selected Qwen3-32B (yang2025qwen3) as its base model. This choice highlights a core advantage of DDR: its flexibility in selecting a base dependency retrieval model without being restricted to specific embedding models.
During the decoding stage for all experiments, the temperature was set to 0.7, and we generated 8 samples for each problem. To ensure a fair comparison, the dependency retrieval step for the DDR method was performed only once, and its results were reused for all subsequent Lean 4 statement formalization generations. It is noteworthy that none of the autoformalization models discussed in the main text, including the fine-tuned ones, have undergone SFT on the FineLeanCorpus (peng2025criticlean).
For a fair comparison, we adopt an experimental setup identical to that of liu2025rethinking. We used the same prompt as described in Section A.7.1 of liu2025rethinking for evaluation on BEq. Similarly, the prompt for the ICL method is also consistent with that in Section A.7.3 of liu2025rethinking. Gemini 2.5 pro is utilized as an auxiliary tool for BEq. Gemini 2.5 pro is also the LLM judge used in section A.2.
Our proposed DDR method is fine-tuned using the Swift (zhao2025swift) framework with the following hyperparameters:
  1. Max Sequence Length: 8192
  2. Batch size: 1
  3. Gradient Accumulation: 64
  4. Training Devices: 8
  5. Train Epochs: 6
  6. Optimizer: AdamW with learning rate $1\times 10^{-4}$, $\beta=(0.9,0.95)$, weight decay 0.1, maximal gradient norm 1.0, no warmup.
  7. Learning Rate Scheduler: Cosine.
  8. LoRA: rank 32, alpha 32, dropout 0.05.
Appendix D Prompt Templates
D.1 Prompt Template for Dependency Retrieval
You are an expert Lean 4 ‘Mathlib‘ dependency extractor.
Your task is to analyze a given mathematical proposition and identify the key mathematical concepts that would require specific definitions, theorems, or notations from the Lean 4 ‘Mathlib‘ library.
**CRITICAL INSTRUCTIONS:**
1. Your output **MUST** be a single, valid JSON object.
2. The JSON object **MUST** have one key: "dependency".
3. The value for the "dependency" key **MUST** be a list of strings. Each string is a potential ‘Mathlib‘ dependency.
4. The dependency names can be fully qualified (e.g., ‘Nat.minFac‘) or common shortcuts (e.g., ‘ncard‘, ‘Icc‘, ‘MOD‘).
5. If there is no potential ‘Mathlib‘ library dependency, output {"dependency":["No usage"]}
6. Do **NOT** write any Lean code. Do **NOT** provide any explanations or text outside of the JSON object. Your entire response must be only the JSON.
Here are some examples of the required input and output:
**Example 1**
**Math Proposition:**
[…]
1. f(n) is an integer for every integer n;
2. 2011^2 $\mid$ f(i) - a_i for i = 1, 2, $\ldots$, 2011;
3. 2011^2 $\mid$ f(n + 2011) - f(n) for every integer n;
is divided by 1000 is equal to 211.
**Output:**
dependency:
[Polynomial,degree,Fin,eval,Set.ncard]
**Example 2**
**Math Proposition:**
Prove that the set of integers (n $\ge$ 1) for which the smallest prime divisor of ((n!)^n + 1) is at most (n + 2015) is finite.
**Output:**
dependency:[Nat.factorial,Finite,
Nat.minFac]
**Example 3**
**Math Proposition:**
[…]
**Output:**
dependency:[MOD,Nat.Prime,Odd,Finset.Icc]
**Example 4**
**Math Proposition:**
Prove that if t represents the number ten, then the value of the expression (t + t + t + t) is 40.
**Output:**
dependency:[No usage]
Now, perform this task for the following proposition.
**Math Proposition:**
{input}
**Output:**
D.2 Prompt Template for Autoformalization with Dependency Retrieval
Please translate mathematical propositions into Lean 4 theorems. ‘Mathlib‘ is the MUST import.
DO NOT try to prove the theorem, ONLY translate it.
Use ‘sorry‘ as placeholder. Use the given potential ‘Mathlib‘ dependency items.
Here are some examples:
Math Proposition:
[…]
[Real,Continuous,Tendsto,Nat,atTop].
Lean Theorem:
import Mathlib
theorem exercise
(f : $\r Real$)
[…]
Math Proposition:
Let G be a finite group (with a multiplicative operation), and A be a subset of G that contains more than half of G’s elements. Prove that every element of G can be expressed as the product of two elements of A. The potential ‘Mathlib‘ dependency are
[Group,Finite,Set,ncard,Nat.card,Rat].
Lean Theorem:
import Mathlib
open Filter Topology
theorem exercise
[Group G]
(hG : Finite G)
(A : Set G)
(hA : A.ncard > (Nat.card G : Rat)/2):
[…]
Math Proposition:
[…]
[Nat,Nat.Prime,ZMod,ncard,MOD].
Lean Theorem:
import Mathlib
theorem exercise
(p : Nat)
(hp : Nat.Prime $\and p$ > 5)
[…]
Math Proposition:
Let p be 10, and q be 15, prove that p+b equals 25. The potential ‘Mathlib‘ dependency are
[No usage].
Lean Theorem:
import Mathlib
theorem exercise
(p : Nat)
(q : Nat)
(hp : p = 10)
(hq : q = 15):
p + q = 25 := sorry
Please translate the following proposition:
Math Proposition:
{input}
The potential ‘Mathlib‘ dependency are
{retrieved-dependency}
Lean Theorem: ‘‘‘
D.3 Other Templates
Other ICL templates employed in this paper, including those for autoformalization without dependency retrieval and BEq verification, adhere to the corresponding settings in liu2025rethinking.